THE ROTATIONAL SPECTRUM OF CIS- AND TRANS-HSSOH

\underline{MONIKA KOERBER}, OLIVER BAUM, THOMAS F. GIESEN, and STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany.

In this talk we present the first pure rotational lines of Oxatrisulfane, HSSOH. Oxatrisulfane is one link between the well-known skew chain molecules HSSSH and HOOOH. The structure of trisulfane HSSSH has been determined using rotational spectroscopy by Liedke et al.^{a} showing HSSSH to form two stable conformers, trans and cis, with the transconformer to be slightly more stable. Only recently hydrogen-trioxide HOOOH has been discovered by Suma et al.^{b} observing the rotational spectra of the trans conformer. For this molecule no evidence for the cis structure has been found. First experimental evidence for the formation of HSSOH was given by Königshofen et al.^{c} via flash vacuum pyrolysis of t-BuS(O)St-Bu at T > 700^{∘}C using IR-spectroscopy to monitor the pyrolysis products. We employed this method to synthesize HSSOH and performed first high-resolution millimeterwave measurements in the range of 75–120 GHz by introducing an all solid state spectrometer. More than 170 lines of the cis conformer and more than 30 lines of trans-HSSOH have been undoubtly identified. Both conformeres are close to the limiting case of a prolate symmetric top with κ_{cis} = −0.925 and κ_{trans} = −0.924 respectively. Their permanent dipole moment points mainly along the b- and cprinciple axes and therefore the spectrum displays a clear perpendicular structure. For analyzing the identified lines a standard S-reduced Hamiltonian has been used and we were able to fit the rotational parameters A, B, C as well as the centrifugal distortion parameters D_{J}, D_{JK}, D_{K}, d_{1} and d_{2}. The c-type transitions of cis-HSSOH are approx. 14 times more intense than the corresponding b-type transitions, whereas in case of trans-HSSOH the intensities of the b- and c-type transitions are almost equal. All experimentally derived parameters are in excellent agreement with those obtained by high level ab initio calculations^{d}.

aLiedtke et al., J. Mol. Struct. 413 265–270 (1997)

bSuma et al., JACS 127 14998–14999 (2005)

cKönigshofen et al., Z. Anorg. Allg. Chem. 625 1779–1786 (1999)

dJ. Gauss, private communication